OpenStudy (anonymous):

An object projected up with initial velocity v attains a height h. Another object projected up with initial velocity 2v will attain what height? I know the answer is 4h but I can't see why.

OpenStudy (calculusxy):

Here is how: h is directly proportional to square of v. so le v0 be v1 and 2v0 be v2 and h1 is h.Here we should find h2. so from the relation h1/h2 = v1^2/v2^2 h/h2 = v0^2/(2v0)^2 h/h2 = 1/4 Therefore h2 = 4h

OpenStudy (anonymous):

Due to the earth's gravity constant (9.8 m/s/s), the object will fall in the shape of a parabola.  If initial velocity was 1: \[1^{2} = 1\] If initial velocity was 2: \[2^{2} = 4\] Hope this helps.